solve dy/dx = x^3(y^2 + 1 )
1
Expert's answer
2020-12-17T15:08:51-0500

Solution: The Given Ordinary Differential Equation is First order separable ODE form. A first order separable ODE has the form "N(y)~y'=M(x)" .

Therefore rewrite in the form of a first order separable form ODE

"\frac{1}{1+y^2}~y'=x^3"

Integrate both sides with respect to "x"

"\int\frac{1}{1+y^2}~y'dx=\int x^3dx"

"\int\frac{1}{1+y^2}~y'dx=\int\frac{1}{1+[y(x)]^2}~y'(x)dx=\int\frac{1}{1+y^2}~\frac{dy}{dx}dx=\int\frac{1}{1+y^2}~dy=\int x^3dx"

"\therefore" considering,


"\int\frac{1}{1+y^2}~dy=\int x^3dx" [using integration formula's]

"arctan(y)=\frac{x^4}{4}+c~~~~or ~~~~~tan^{-1}y=\frac{x^4}{4}+c"


"\therefore y=tan(\frac{x^4}{4}+c)"

This is general solution of the given differential equation.